bHelp needed with missing number:
Select the missing number
9, 21, 35, ?, 69
Solution:
a)50
b)51
c)60
d)61

Answer by farohw(175)


The pattern is add 12, 14, 16, 18 to each number.

Thus, 9 + 12 = 21 and 21 + 14 = 35 and 35 + 16 = 51 and 51 + 18 = 69.

So your answer is b) 51